OpenStudy (anonymous):

let n be the middle number of three consecutive integers. write an expression for the sum of these integers.

OpenStudy (anonymous):

Hm...I'm trying to think of a way to give you a hint without explicitly telling you the solution. For example, 1,2, and 3 are consecutive integers. If 2 is the middle integer in this sequence, how can you express 2 in terms of the first and the third integer? :-)

OpenStudy (anonymous):

I have no idea lol this is new to me

OpenStudy (anonymous):

Considering the example 1,2,3 again, if 2 is the middle number, what do I do to 2 in order to get 1? What do I do to 2 in order to get 3? :-)

OpenStudy (anonymous):

(It's easier than it looks)

OpenStudy (anonymous):

in order to get 3 u add so it'd be like n+2 and to get 1 u would subract so 2-1

OpenStudy (anonymous):

I think you mean 2+1. ;-) Anyways, then it follows that we can rewrite the consecutive integers 1,2,3 in terms of the middle integer! In particular, 1,2,3 is the same as 2-1,2,2+1. Once you have this, it's very easy to generalize the result. In this case, the problem is asking for the 3 consecutive integers given that n is the second integer. Do you think you can come up with the appropriate consecutive integers now? :-)

OpenStudy (anonymous):

is it (n-1)+n+(n+1)?

OpenStudy (anonymous):

Yes it is. :-) Do you know if they want you to simplify it?

OpenStudy (anonymous):

yeah or you could just leave the problem as it is

OpenStudy (anonymous):

Generally, I think it would be best to simplify it. What do you get in this case?

OpenStudy (anonymous):

3n

OpenStudy (anonymous):

Yup, that's it. :-) Does this make a little more sense now? :-)

OpenStudy (anonymous):

Yes it does, thanks :) can you answer another question for me?

OpenStudy (anonymous):

Sure. :-)

OpenStudy (anonymous):

The sum of the four angle measures of any convex quadrilateral is 360 degrees. Suppose that a convex quadrilateral has angle measures of 30 , y+10 , and 10y-8 degrees. Write an expression for the degree measure of the fourth angle.

OpenStudy (anonymous):

Let the 4th degree measure be \(x\). Then by what you're told in the problem, we have that \(30+(y+10)+(10y-8)+x = 360\). What do you get when you solve this equation for \(x\)? :-)

OpenStudy (anonymous):

Do we distribute?

OpenStudy (anonymous):

Well, we're not multiplying anything together, so there's no need to distribute. What you want to do though is combine like terms and solve for \(x\).

OpenStudy (anonymous):

Also, because you're adding everything together, you can drop the parenthesis. I just included them to emphasize the different angles that were being added together. :-)

OpenStudy (anonymous):

okay so is it gonna be 58+11y=360?

OpenStudy (anonymous):

48+11y=360

OpenStudy (anonymous):

omg im so sorry lol its 5y+10

OpenStudy (anonymous):

30,5y+10 and 10y-8

OpenStudy (anonymous):

so it will be 48+15y=360

OpenStudy (anonymous):

We're trying to solve for x. When you combine everything on the LHS, we see that \(30+y+10+10y-8 +x = x + 11y + 32\). If that's equal to 360, then \(x=360 - (11y+32) = \ldots\). Can you take things from here?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

I got 328-11y

OpenStudy (anonymous):

Yes, that's correct! :-)

OpenStudy (anonymous):

do we divide now or is that the answer?

OpenStudy (anonymous):

No, that's our answer. The 4th angle measurement in terms of the 3 that were given is \(328-11y\). Does this make more sense than it did originally? :-)

OpenStudy (anonymous):

Yes it makes alot more sense haha

OpenStudy (jacob902):

Chance, don't think too much! If the middle number is n, then the first number is n-1 and the third number is n + 1. So, we have (n - 1) + n + (n + 1). In this case, you can take out the parentheses. Combine like terms and you get 3n. Let's check. 3 + 4 + 5 = 12 = 3(4) 5 + 6 + 7 = 18 = 3(6). Hope this helps!